chances of solving a probability problem indepednently by four students a,b,c,d are 1/2,1/4,1/6 and 1/8.if all of them try to solve the problem .what is the probability that the problem will be solved by at least by one of them
1
Expert's answer
2012-09-06T11:03:51-0400
Probability for all of them being wrong is P=1/2*(1-1/4)*(1-1/6)*(1-1/8)=0.091.
Now probability we're looking for is as follows 1-0.091=0.909